Given the coordinates of the vertices of a quadrilateral, determine whether it is a square, a rectangle or a
parallelogram. Then find the area of the quadrilateral.
J(–6, 4), K(–1, 4), L(–6, –4), M(–1, –4)
Thank you!!

Answer by ikleyn(52776)

.
The side  JK  is horizontal y = 4. Its length is -1 - (-6) = 5 units.


The side  LM is horizontal y = -4. Its length is -1 - (-6) = 5 units.


The side  JL  is vertical y = x= -6.  Its length is 4 - (-4) = 8 units.


The side  KM  is vertical y = x= -1.  Its length is 4 - (-4) = 8 units.


The quadrilateral  JKML  (! in this order !)  is a rectangle with dimensions 5 and 8 units.


Its area is 5*8 = 40 square units.